please show me how to get this equation. write the equation of the line which passes through the point(2,1) and is parallelto the line with equation y=5/2x+6 Thank you.
Answer by checkley71(8403)

Y=5/2X+6 HAS A SLOPE OF 5/2 AND PASSES THROUGH POINTS(2,1) THEN
1=5/2*2+b
1=5+b
B=1-5
b=-4 THUS THE EQUATION FOR THIS LINE IS
Y=5/2X-4